# High priority targets for transient gravitational waves from glitching pulsars

### v1

@@ Abstract
Glitching pulsars are expected to be important sources of gravitational waves. In this paper, we explore six different models that propose the emission of transient continuous waves, lasting days to months, coincident with glitches. The maximal gravitational wave energy is calculated for each model, which is then used to determine whether associated gravitational waves could be detectable with LIGO-Virgo-KAGRA’s O4 detectors. We provide an analytical approximation to calculate the signal-to-noise ratio which includes information about the source’s sky position, improving on previous estimates that assume isotropic or sky and orientation averaged sensitivities. Applying the calculation to the entire glitching population, we find that certain models predict detectable signals in O4, whereas others do not. We also rank glitching pulsars in order of how significant a signal would be, based on archival data, and we find that for all models, the Vela pulsar (PSR J0835--4510) would provide the strongest signal. Moreover, PSR J0537--6910 is not expected to yield a detectable signal in O4, but will start becoming relevant for next generation detectors. Our analysis also extends to the entire pulsar population, regardless of whether they have glitched or not, and we provide a list of pulsars that would present a significant signal, if they were to glitch. Finally, we apply our analysis to the latest April 2024 Vela glitch and find that a signal should be detectable under certain models. The non-detection of a supposedly detectable signal would provide an efficiency factor that quantifies how much a model can contribute to gravitational wave emission, eventually leading to a differentiation of models and independent constraints on physical parameters.
@@ 1 Introduction
The other well-known case, which will be the case that we will focus on here, is from NS glitches. A glitch refers to the sudden increase in the spin of a NS, thought to be due to the unpinning of superfluid vortices (Anderson & Itoh, 1975). This unpinning causes a rapid transfer of angular momentum from the fast superfluid interior to the crust, which the magnetosphere and electromagnetic radiation is tied to. There is also another popular model, though disfavoured for NSs with large and frequent glitches, which is called the starquake model (Ruderman, 1969; Baym & Pines, 1971). This model attributes glitches to a rapid decrease in the moment of inertia of the NS when the crust’s elastic limit is exceeded during the spin-down of a NS. A review of these glitch models and others can be found in Haskell & Melatos (2015). For some glitches, there is also a glitch recovery where the spin recovers back to, but not necessarily reaching, the pre-glitch spin frequency. This process normally takes days to months (Antonopoulou et al., 2022; Zhou et al., 2022), similar to the timescales of interest that define transient CWs.
@@ 2.1.1 Starquake model (one component)
As the NS spins-down (primarily through magnetic dipole radiation), the centrifugal force becomes weaker, leading to the NS wanting to become more spherical. However, due to the rigidity of the crust, this is prevented from happening at the expense of strain being built up.
@@ 2.1.2 Vortex unpinning model (two components)
We will now move onto a more realistic two component set-up for describing NSs. There is great reason to consider NSs with (at least) two components with (at least) one of them being superfluid; without superfluidity, there are several observed phenomena that are difficult to explain such as: large and frequent glitches (Baym & Pines, 1971; Pines & Alpar, 1985), long post-glitch recovery timescales (Baym et al., 1969a, b; Alpar et al., 1984; Pines & Alpar, 1985), and long thermal cooling timescales (Shibazaki & Lamb, 1989; Page & Applegate, 1992; Larson & Link, 1999). More recently, observations of glitches with a “delayed spin-up” (Shaw et al., 2018; Ashton et al., 2019; Shaw et al., 2021) give observational evidence for there being three components (Graber et al., 2018), but we will not cover this here and leave it for future work (Yim & Jones, in prep.). For our purposes, two components are enough as this is all that is needed to explain glitch rises and post-glitch recoveries (Anderson & Itoh, 1975; Alpar et al., 1984).
@@
As the NS spins-down more, the lag grows and it becomes harder for the vortices to remain pinned. At some critical lag, the vortices collectively unpin and rapidly transfer their angular momentum to the unpinned component which is what we observe as a glitch. Consequently, the pinned component spins-down and the unpinned component spins-up, resulting in the lag decreasing. Usually, the lag is assumed to be zero after the glitch so that both components co-rotate temporarily.
@@ 2.2.1 Transient mountain model
If one uses $Q=1/2$, we get the same energy as the starquake model and for $Q=1$, we get the same as the “naïve” model that will be covered later in Section 2.3.1, where only energy conservation is considered. The value of $Q$ varies for different pulsars and even for different glitches from the same pulsar, but to get an idea of their values, the average $Q$ for the Crab and Vela pulsars are around $\langle Q\rangle\sim 0.8$ and $\langle Q\rangle\sim 0.2$, respectively (Crawford & Demiański, 2003; Yim & Jones, 2020).
@@ 3.2 Analytic approximations
The explicit form of $A_{2}$ is provided in Appendix A. Eq. (40) is what we propose as a suitable choice for calculating the SNR of transient CWs, given $T_{\text{GW}}$ is sufficiently large. Moreover, this choice of $\beta$ allows for the inclusion of additional information, like the declination of the source, polarisation angle, inclination, detector latitude and detector orientation. Note that when using this approximation, the right ascension is not important. This makes sense as for a sufficiently long signal, the Earth would rotate about its axis at least once, meaning each value of the right ascension would carry equal weighting.
@@ 3.3 Accuracy of approximation
As it may be important for future observations, we also provide figures similar to Fig. 2 but specifically for three pulsars of interest: the Crab pulsar, the Vela pulsar and PSR J0537--6910 (Antonopoulou et al., 2018; Ferdman et al., 2018; Ho et al., 2020b, 2022; Fesik & Papa, 2020; Abbott et al., 2021a, b). These are shown in Figs. 5 - 7. One finds that the approximation, eq. (40), gives an error smaller than 10% so long as $T_{\text{GW}}\geq 0.36~{}\text{d},0.77~{}\text{d},0.30~{}\text{d}$ for the Crab pulsar, Vela pulsar and PSR J0537--6910, respectively. As for the values of $A_{2}$ for these three pulsars, for $\iota=0\text{\textdegree}$, we have $A_{2}=0.355,0.418,0.365$ for the Crab pulsar, $A_{2}=0.456,0.379,0.432$ for the Vela pulsar and $A_{2}=0.518,0.354,0.498$ for PSR J0537--6910, where each number in each list represents LIGO Hanford, LIGO Livingston and Virgo, respectively.
@@ 4.1 Including glitch data
We present the results of the SNR calculation as histograms in Fig. 8. When using $\beta=A_{2}$, each GW detector has a slightly different SNR due to the dependence of $A_{2}$ on $(\lambda,\gamma)$. Therefore, we choose to provide histograms for the largest SNR obtained from the three detectors. The sensitivities that we use are O4 sensitivities and we obtain the sensitivity curves from LIGO Document T2200043-v3.666https://dcc.ligo.org/LIGO-T2200043/public
@@
We will now look at the most favourable pulsars that we expect to yield a detection. In Tables 3 - 5, we present the top 15 pulsars that give that largest GW energies and SNRs assuming different energy models using $\beta=A_{2}(\iota=0\text{\textdegree})$. For each pulsar, we also present its right ascension, declination, spin frequency, spin frequency derivative, distance, number of observed glitches, glitch size corresponding to the largest SNR, fractional change in the spin-down rate corresponding to the largest SNR, and for the transient mountain model, the healing parameter corresponding to the largest SNR.
@@
It is also worth looking at individual pulsars too, particularly the ones that have already been targeted by LVK. These are the Crab pulsar (PSR J0534+2200), the Vela pulsar (PSR J0835--4510) and the “big glitcher” (PSR J0537--6910). Firstly, we track the Crab pulsar. It ranks 8th, 15th and 3rd in the agnostic, vortex unpinning and transient mountain models, respectively. The drop in ranking for the vortex unpinning model is due to the relatively small glitch sizes that the Crab exhibits, as the GW energy in this model is quadratic in the glitch size rather than linear. The Crab ranks relatively highly for the transient mountain model due to its large $Q$. Moving onto Vela, it remains the highest priority target for all models primarily due to its large glitches and close distance, holding also true for PSR J0940--5428.
@@ 4.2 Independent of glitch data
From here, we see that the Vela pulsar is the pulsar with the 8th highest $\chi$, with a value of $\chi\approx 3.99\times 10^{-20}~{}\text{erg}^{-\frac{1}{2}}$. We also report the values for the Crab pulsar and PSR J0537--6910 which are $\chi\approx 5.61\times 10^{-21}~{}\text{erg}^{-\frac{1}{2}}$ and $\chi\approx 2.60\times 10^{-22}~{}\text{erg}^{-\frac{1}{2}}$, respectively. We can verify our earlier calculations, for example, for Vela, by substituting $E_{\text{GW}}=1.53\times 10^{43}~{}\text{erg}$, $2.35\times 10^{39}~{}\text{erg}$ and $1.50\times 10^{42}~{}\text{erg}$ into eq. (48) for the agnostic, vortex unpinning and transient mountain models and we recover the maximum SNRs in Tables 3 - 5 to within some rounding error. This makes $\chi$ a powerful tool to quickly estimate the SNR if $E_{\text{GW}}$ is provided. Additionally, the calculation no longer needs to be restricted to glitches but could also be used for other transient phenomena such as magnetar bursts or fast radio bursts, if one assumes $f=2\nu$ radiation is emitted.
@@ 6 Discussion
If $\iota$ is non-zero, then one might ask what a suitable value would be. We can obtain this from electromagnetic observations, specifically from radio polarisation data, which are often fit with the “rotating vector model” (Radhakrishnan & Cooke, 1969; Everett & Weisberg, 2001). Using this model, one obtains a value of the magnetic inclination angle ($\alpha_{\text{RVM}}$, the angle between the pulsar’s rotation axis and magnetic axis) and the viewing angle ($\beta_{\text{RVM}}$, the angle between the magnetic axis and the observer’s line of sight). The inclination $\iota$ is therefore just the sum of these two angles, $\iota=\alpha_{\text{RVM}}+\beta_{\text{RVM}}$ (or $\iota=180\text{\textdegree}-\alpha_{\text{RVM}}-\beta_{\text{RVM}}$, see Jones (2015) and Abbott et al. (2017)). Future work should incorporate this inclination dependence into the analysis to get a more faithful SNR calculation.
@@ 7 Conclusion
Next, we no longer restricted ourselves to the population of glitching pulsars and applied the analysis to the entire pulsar population, assuming that one day, some of these pulsars may glitch. We provided two lists of high priority sources (Tables 6 and 7), and in the latter, the top 164 pulsars have not been observed to glitch before. We advocate for better timing programmes for these potentially interesting pulsars.

### v2

@@ Abstract
Glitching pulsars are expected to be important sources of gravitational waves. In this paper, we explore six different models that propose the emission of transient continuous waves, lasting days to months, coincident with glitches. The maximal gravitational wave energy is calculated for each model, which is then used to determine whether associated gravitational waves could be detectable with LIGO-Virgo-KAGRA’s O4 detectors. We provide an analytical approximation to calculate the signal-to-noise ratio which includes information about the source’s sky position, improving on previous estimates that assume isotropic or sky and orientation averaged sensitivities. By analysing the entire glitching population, we find that certain models predict detectable signals in O4, whereas others do not. We also rank glitching pulsars by signal-to-noise ratio, based on archival data, and we find that for all models, the Vela pulsar (PSR J0835--4510) would provide the strongest signal. Moreover, PSR J0537--6910 is not expected to yield a detectable signal in O4, but will start becoming relevant for next generation detectors. Our analysis also extends to the entire pulsar population, regardless of whether they have glitched, and we provide a list of pulsars that would present a significant signal, if they were to glitch. Finally, we apply our analysis to the April 2024 Vela glitch and find that a signal should be detectable under certain models. The non-detection of a supposedly detectable signal would provide an efficiency factor that quantifies a model’s contribution to gravitational wave emission, eventually leading to a differentiation of models and independent constraints on physical parameters.
@@ 1 Introduction
The other well-known case, which will be the case that we will focus on here, is from NS glitches. A glitch refers to the sudden increase in the spin of a NS, thought to be due to the unpinning of superfluid vortices (Anderson & Itoh, 1975). This unpinning causes a rapid transfer of angular momentum from the fast superfluid interior to the crust, which the magnetosphere and electromagnetic radiation are tied to. There is also another popular model, though disfavoured for NSs with large and frequent glitches, which is called the starquake model (Ruderman, 1969; Baym & Pines, 1971). This model attributes glitches to a rapid decrease in the moment of inertia of the NS when the crust’s elastic limit is exceeded during the spin-down of a NS. A review of these glitch models and others can be found in Haskell & Melatos (2015). For some glitches, there is also a glitch recovery where the spin recovers back to, but not necessarily reaching, the pre-glitch spin frequency. This process normally takes days to months (Antonopoulou et al., 2022; Zhou et al., 2022), similar to the timescales of interest that define transient CWs.
@@ 2.1.1 Starquake model (one component)
As the NS spins-down (primarily through magnetic dipole radiation), the centrifugal force becomes weaker, leading to the NS wanting to become more spherical. However, due to the rigidity of the crust, this is prevented from happening at the expense of elastic strain being built up.
@@ 2.1.2 Vortex unpinning model (two components)
We will now move onto a more realistic two component set-up for describing NSs. There is great reason to consider NSs with (at least) two components with (at least) one of them being superfluid; without superfluidity, there are several observed phenomena that are difficult to explain such as: large and frequent glitches (Baym & Pines, 1971; Pines & Alpar, 1985), long post-glitch recovery timescales (Baym et al., 1969a; Baym et al., 1969b; Alpar et al., 1984; Pines & Alpar, 1985), and long thermal cooling timescales (Shibazaki & Lamb, 1989; Page & Applegate, 1992; Larson & Link, 1999). More recently, observations of glitches with a “delayed spin-up” (Shaw et al., 2018; Ashton et al., 2019; Shaw et al., 2021) give observational evidence for there being three components (Graber et al., 2018), [^1] See also Haskell et al. (2018) who were able to use a two-component model to explain delayed spin-ups. The key ingredient that allowed them to do this was to have vortex accumulation and differential rotation, which affected the coupling of different parts of the star (Khomenko & Haskell, 2018). but we will not cover this here and leave it for future work (Yim & Jones, in prep.). For our purposes, two components are enough as this is all that is needed to explain glitch rises and post-glitch recoveries (Anderson & Itoh, 1975; Alpar et al., 1984).
@@
As the NS spins-down more, the lag grows and it becomes harder for the vortices to remain pinned. At some critical lag, the vortices collectively unpin and rapidly transfer their angular momentum to the unpinned component which is what we observe as a glitch. Consequently, the pinned component spins-down and the unpinned component spins-up, resulting in the lag decreasing. Usually, the lag is assumed to be zero after the glitch so that both components co-rotate temporarily. However, this assumption is still unclear and model-dependent. For simplicity, we will assume co-rotation immediately after the glitch.
@@ 2.2.1 Transient mountain model
If one uses $Q=1/2$, we get the same energy as the starquake model [^2] This is only to recover the same energy and not a prediction of the starquake model in general. and for $Q=1$, we get the same as the “naïve” model that will be covered later in Section 2.3.1, where only energy conservation is considered. The value of $Q$ varies for different pulsars and even for different glitches from the same pulsar, but to get an idea of their values, the average $Q$ for the Crab and Vela pulsars are around $\langle Q\rangle\sim 0.8$ and $\langle Q\rangle\sim 0.2$, respectively (Crawford & Demiański, 2003; Yim & Jones, 2020).
@@ 3.2 Analytic approximations
The explicit form of $A_{2}$ is provided in Appendix A. Eq. (40) is what we propose as a suitable choice for calculating the SNR of transient CWs, given $T_{\text{GW}}$ is sufficiently large. Moreover, this choice of $\beta$ allows for the inclusion of additional information, like the declination of the source, polarisation angle, inclination, detector latitude and detector orientation. Note that when using this approximation, the right ascension is not important. This makes sense as for a sufficiently long signal, the Earth would rotate about its axis at least once, meaning each value of the right ascension would carry equal weighting. Table 2 summarises the different options for $\beta$, their respective assumptions as well as where each choice is applicable.
Table 2: A table containing the different options for $\beta$, their respective assumptions and the scenarios where their use may be appropriate.
$\beta$ Assumptions When applicable?
1 Most optimistic Order-of-magnitude calculations or short GW signals
$\frac{4}{25}$ Sky position and orientation averaged over Same scenarios as $\beta=1$ case
$A_{2}$ Sufficiently long transient signal Have sky position and/or inclination angle of the source
@@ 3.3 Accuracy of approximation
As it may be important for future observations, we also provide figures similar to Fig. 2 but specifically for three pulsars of interest: the Crab pulsar, the Vela pulsar and PSR J0537--6910 (Antonopoulou et al., 2018; Ferdman et al., 2018; Ho et al., 2020b; Ho et al., 2022; Fesik & Papa, 2020; Abbott et al., 2021a, b). These are shown in Figs. 5 - 7. One finds that the approximation, eq. (40), gives an error smaller than 10% so long as $T_{\text{GW}}\geq 0.36~{}\text{d},0.77~{}\text{d},0.30~{}\text{d}$ for the Crab pulsar, Vela pulsar and PSR J0537--6910, respectively. As for the values of $A_{2}$ for these three pulsars, for $\iota=0\text{\textdegree}$, we have $A_{2}=0.355,0.418,0.365$ for the Crab pulsar, $A_{2}=0.456,0.379,0.432$ for the Vela pulsar and $A_{2}=0.518,0.354,0.498$ for PSR J0537--6910, where each number in each list represents LIGO Hanford, LIGO Livingston and Virgo, respectively.
@@ 4.1 Including glitch data
We present the results of the SNR calculation as histograms in Fig. 8. When using $\beta=A_{2}$, each GW detector has a slightly different SNR due to the dependence of $A_{2}$ on $(\lambda,\gamma)$. Therefore, we choose to provide histograms for the largest SNR obtained from the three detectors. The sensitivities that we use are O4 sensitivities and we obtain the sensitivity curves from LIGO Document T2200043-v3.888https://dcc.ligo.org/LIGO-T2200043/public
@@
We will now look at the most favourable pulsars that we expect to yield a detection. In Tables 4 - 6, we present the top 15 pulsars that give that largest GW energies and SNRs assuming different energy models using $\beta=A_{2}(\iota=0\text{\textdegree})$. For each pulsar, we also present its right ascension, declination, spin frequency, spin frequency derivative, distance, number of observed glitches, glitch size corresponding to the largest SNR, fractional change in the spin-down rate corresponding to the largest SNR, and for the transient mountain model, the healing parameter corresponding to the largest SNR.
@@
It is also worth looking at individual pulsars too, particularly the ones that have already been targeted by LVK. These are the Crab pulsar (PSR J0534+2200), the Vela pulsar (PSR J0835--4510) and the “big glitcher” (PSR J0537--6910), of which the latter has been suggested to also emit GWs via $r$-modes (e.g. Fesik & Papa, 2020; Abbott et al., 2021b). Firstly, we track the Crab pulsar. It ranks 8th, 15th and 3rd in the agnostic, vortex unpinning and transient mountain models, respectively. The drop in ranking for the vortex unpinning model is due to the relatively small glitch sizes that the Crab exhibits, as the GW energy in this model is quadratic in the glitch size rather than linear. The Crab ranks relatively highly for the transient mountain model due to its large $Q$. Moving onto Vela, it remains the highest priority target for all models primarily due to its large glitches and close distance, holding also true for PSR J0940--5428.
@@ 4.2 Ranking all pulsars
From here, we see that the Vela pulsar is the pulsar with the 8th highest $\chi$, with a value of $\chi\approx 3.99\times 10^{-20}~{}\text{erg}^{-\frac{1}{2}}$. We also report the values for the Crab pulsar and PSR J0537--6910 which are $\chi\approx 5.61\times 10^{-21}~{}\text{erg}^{-\frac{1}{2}}$ and $\chi\approx 2.60\times 10^{-22}~{}\text{erg}^{-\frac{1}{2}}$, respectively. We can verify our earlier calculations, for example, for Vela, by substituting $E_{\text{GW}}=1.53\times 10^{43}~{}\text{erg}$, $2.35\times 10^{39}~{}\text{erg}$ and $1.50\times 10^{42}~{}\text{erg}$ into eq. (48) for the agnostic, vortex unpinning and transient mountain models and we recover the maximum SNRs in Tables 4 - 6 to within some rounding error. This makes $\chi$ a powerful tool to quickly estimate the SNR if $E_{\text{GW}}$ is provided. Additionally, the calculation no longer needs to be restricted to glitches but could also be used for other transient phenomena such as magnetar bursts or fast radio bursts, if one assumes $f=2\nu$ radiation is emitted.
@@ 6 Discussion
If $\iota$ is non-zero, then one might ask what a suitable value would be. We can obtain this from electromagnetic observations, specifically from radio polarisation data, which are often fit with the “rotating vector model” (Radhakrishnan & Cooke, 1969; Everett & Weisberg, 2001; Jones, 2007). Using this model, one obtains a value of the magnetic inclination angle ($\alpha_{\text{RVM}}$, the angle between the pulsar’s rotation axis and magnetic axis) and the viewing angle ($\beta_{\text{RVM}}$, the angle between the magnetic axis and the observer’s line of sight). The inclination $\iota$ is therefore just the sum of these two angles, $\iota=\alpha_{\text{RVM}}+\beta_{\text{RVM}}$ (or $\iota=180\text{\textdegree}-\alpha_{\text{RVM}}-\beta_{\text{RVM}}$, see Jones (2015) and Abbott et al. (2017)). Future work should incorporate this inclination dependence into the analysis to get a more faithful SNR calculation.
@@ 7 Conclusion
Next, we no longer restricted ourselves to the population of glitching pulsars and applied the analysis to the entire pulsar population, assuming that one day, some of these pulsars may glitch. We provided two lists of high priority sources (Tables 7 and 8), and in the latter, the top 164 pulsars have not been observed to glitch before. We advocate for better timing programmes for these potentially interesting pulsars.